# Still More Shades of Null: An Evaluation Suite for Responsible Missing Value Imputation

### v3

@@ 5.3. Stability of the Predictive Model
Training set missingness. Figure 10 shows Label Stability of the Random Forest model on diabetes as a function of training set missingness rate. Under MCAR and MAR missingness, most MVI techniques with the exception of boostclean, deletion, and clustering show good stability (comparable to the model trained on clean data), and are generally insensitive to missingness rates. deletion and clustering are the least stable methods and show a monotonic decrease in stability with increase in missingness rate, with strongest effect under MNAR. auto-ml and miss-forest are the most robust MVI techniques and preserve stability of the clean model under all missingness settings and error rates.
@@
Figure 47 in Appendix 46 shows the Label Stability of the Random Forest model on diabetes as a function of test missingness rate. We find that missingness rate has little effect on model stability under all missingness settings, with the exception of clustering, which shows lower label stability at higher missingness rates, most pronounced in scenario S7 (MAR train, MNAR test).
@@ 6. Summary of Experimental Findings
Do not drop your nulls! Building on evidence in the scientific community (Little and Rubin, 1987; Martínez-Plumed et al., 2019; Little et al., 2015; Joel and Doorsamy, 2022), our findings indicate that, in addition to hurting predictive performance, deletion also worsens stability and fairness, arguing against the use of this method in practice.
@@
There is conflicting evidence in the literature on the performance of multiple imputation (MI) (Joost R. van Ginkel and van der Voort, 2020), and our empirical findings are also mixed and somewhat unexpected.  Feng (2023) and  Le Morvan et al. (2021) argue that MI has superior predictive performance compared to impute-then-classify approaches.  Graham (2009) and McNeish (2017) find that MI performs well even with less data and small error rates. In contrast, we find that MI (specifically, boostclean) is only competitive on small datasets, and is unstable compared to simpler MVI techniques. This can likely be explained by the complexity-stability trade-offs (Domingos, 2000): MI uses a more complex model class and is able to minimize the empirical loss, but due to this very complexity it shows a large variation in predictions under small train set perturbations.
@@
Shadbahr et al. (2023) posit that distributional imputation quality metrics are a more reliable indicator of a model’s predictive performance than discrepancy metrics. We find that neither discrepancy nor distributional metrics are sufficiently indicative of downstream model performance (according to F1), indicating that poor imputation quality can be overcome by a strong learner. Further, we find that imputation fairness metrics are not predictive of model fairness: perfect imputation fairness can result in downstream model unfairness. Conversely, models trained using MVI techniques with imputation unfairness can still have good downstream fairness.
@@
Figure 11. Spearman correlation ($\rho$) between MVI technique, model type, and performance metrics (F1, fairness and stability) for different train missingness mechanisms (subplots). TPRD and TNRD values close to 0 are ideal (fair), so we compute correlations using $TPRD\_Reversed=1-|TPRD|$ and $TNRD\_Reversed=1-|TNRD|$. Supplementary plots for each combination of training and test missingness are in Appendix E.
@@
The predictive performance of models (according to F1) is more sensitive to missingness rates during testing than during training. Fairness, on the other hand, is highly sensitive to both training and test missingness rates. Lastly, there is little effect of test missingness rates on model stability, but most MVI techniques are more unstable at higher training missingness rates.
No good methods for MNAR missingness.
MNAR is theoretically the hardest setting to model. Empirically, we find that MVI techniques perform poorly with MNAR during testing. Conversely, even methods that are trained on MNAR but evaluated on “simpler” test missingness (MCAR and MAR) perform poorly. Further, we find that model F1 and stability are more sensitive to missingness rates under MNAR than under MCAR and MAR. Designing MVI techniques that are effective under MNAR is an open research direction.
@@
There has been an explosion of interest in deep learning for tabular data (Hwang and Song, 2023; Gorishniy et al., 2021; Borisov et al., 2022; Ye et al., 2024; Fayaz et al., 2022). Recall that in Figures 3, 4, and 5 we reported results only for the best-performing models (in terms of F1), to streamline exposition. We find that tree-based methods like random forest and gradient boosted trees still outperform new deep table-learning models like GANDALF (Joseph and Raj, 2022) on small datasets (such as diabetes and german), corroborating the findings of Grinsztajn et al. (2022) and Shwartz-Ziv and Armon (2022), on tabular datasets with (simulated) missingness. GANDALF, however, is the best-performing model on larger datasets (heart and folk-employment in our study), underscoring the growing effectiveness of deep learning on a variety of data modalities.
@@
A key novelty in our study is the evaluation under mixed or multi-mechanism missingness— which is a more complex but realistic missingness pattern, likely to be encountered in practice. We find that trends are different under single- and multi-mechanism missingness. Figure 11 shows that evaluation metrics and model types such as gandalf__\__clf and lgbm__\__clf are uncorrelated under single-mechanism missingness, but correlated under mixed missingness, for example for F1 ($\rho$=-0.24 for gandalf__\__clf and $\rho$=-0.37 for lgbm__\__clf) and fairness according to TPRD ($\rho$=0.55 for gandalf__\__clf and $\rho$=-0.4 for lgbm__\__clf).
@@
Prior experimental work (Hasan et al., 2021; Lin and Tsai, 2020; Joel and Doorsamy, 2022; Shadbahr et al., 2023; García Laencina et al., 2010) concluded that there is no definitive answer to the question of which missing value imputation technique is best according to predictive performance. Our findings support this and demonstrate trade-offs between F1, fairness and stability, based on the architecture of the predictive model. For example, from Figure 11, under multi-mechanism missingness, logistic regression (lr__\__clf) is positively correlated with F1 ($\rho$=0.23) and stability ($\rho$=0.38) but negatively correlated with fairness according to TPRD ($\rho$=-0.23). Random forest (rf__\__clf), on the other hand, is weakly positively correlated with fairness according to TPRD ($\rho$=0.16) and F1 ($\rho$=0.25), but negatively correlated with stability ($\rho$=-0.24). Lastly, deep table-learning (gandalf__\__clf) is strongly positively correlated with fairness according to TPRD ($\rho$=0.55), weakly correlated with stability ($\rho$=0.12), and negatively correlated with F1 ($\rho$=-0.24). This is not bad news, but rather an acknowledgement of the inherent complexity of the task of learning from incomplete data, and a call for more rigorous and holistic evaluation protocols— of the kind demonstrated in this work — to select the most suitable missing value imputation technique and model architecture for a given task.
@@ 7. Conclusions, Limitations and Future Work
Conclusions. We presented Shades-of-Null, an evaluation suite for responsible missing value imputation. A key novelty of our work is in evaluating the predictive performance of models trained on imputed data alongside fairness and stability. Another novelty is our modeling of realistic missingness scenarios that go beyond Rubin’s classic settings of MCAR, MAR and MNAR to include multi-mechanism missingness and missingness shift. We showcased results of an extensive study, with 23,940 experimental pipelines, that uses our framework to evaluate a variety of MVI methods in realistic missingness scenarios. We synthesized interesting trends, and identified trade-offs between evaluation metrics.
@@
It is interesting future work to create research datasets with naturally-occurring missingness shifts. Further, while predictive performance, fairness and stability can be seen as orthogonal dimensions of performance, we have uncovered trade-offs among them that arise due to missingness in the data, the choice of a missing value imputation method, and, in some cases, the architecture of the model trained on the data post-imputation. This, coupled with the finding that there is no universally best-performing imputation method, motivates the development of holistic metrics, evaluation procedures, and imputation techniques that are grounded in the lifecycle of data collection, processing and use. That is, much exciting work remains in responsible data engineering, of which missing value imputation is an important dimension, and to which our work makes a contribution.
@@
This supplemental material provides significantly more detail regarding the results presented in this paper. The organization is as follows: We first provide additional details on the implementation of the evaluation suite, which include a broad description of the MVI techniques used in our study (Appendix A.1) and an explanation of the corrections and enhancements made to existing benchmarks and MVI techniques to ensure the development of our truly fair evaluation suite (Appendix A.2). Next, we extend the description of our experimental settings to facilitate reproducibility (Appendix B). Specifically, Appendix B.1 provides detailed information about our experimental setup on the cluster, while Appendix B.2 extends our methodology for simulating missingness in other datasets, including proportions and base rates of protected groups, realistic missingness scenarios, feature correlation coefficients with the target, and feature importance for each dataset. In Appendix C, we offer additional results to those discussed in Section 4. In particular, Appendix LABEL:apdx:single-mechanism-imp-quality presents the imputation quality of all MVI techniques across all datasets from the perspectives of accuracy, distributional alignment, and fairness, while Appendix C.1 showcases the fairness of predictive models measured by metrics beyond the TPRD mentioned in Section 4. Finally, Appendix D presents supplementary experimental findings on the missingness shift discussed in Section 5. Specifically, Appendix D.1 examines the accuracy, stability, and fairness dimensions of model performance under missingness shift with fixed train and test error rates, highlighting its effects on various model types and datasets. Furthermore, Appendix D.2 explores further experimental results on model performance under missingness shift with variable train and test error rates.
@@ A.1. Missing Value Imputation Techniques
In our evaluation suite, we evaluate 8 MVI techniques, classified into 5 main categories based on the taxonomy presented in (Emmanuel et al., 2021). In each category, we selected one or two SOTA methods for our comparison. Each technique is capable of imputing both categorical and numerical features. All configurations and hyperparameters for the MVI techniques described below are specified in the configuration module within the source code.
@@
    Statistical techniques:
    1. (a)
      median-mode: We impute missing values with the median of the complete case for numerical columns and mode (most frequently occurring value) for categorical columns.
    2. (b)
      median-dummy: We impute missing values with the median of the complete case for numerical columns and assign a (new) dummy category for nulls in categorical columns.
@@
    Machine learning approaches:
    1. (a)
      miss-forest (supervised) (Stekhoven and Bühlmann, 2011): This approach iteratively trains a RF model on a set of clean samples (no missingness) and predicts the missing values.
    2. (b)
      clustering (unsupervised) (Gajawada and Toshniwal, 2012): We assign the missing value to a cluster, based on the distance from the cluster center, and then impute the missing value with the mean or mode of the data points in that cluster. [^8] We used the k-prototypes implementation from https://github.com/nicodv/kmodes We implemented this method ourselves, as no existing solution supports null imputation based on clustering for data with both numerical and categorical features.
@@
    Deep learning approaches:
    1. (a)
      datawig (Biessmann et al., 2019): This method implements deep learning modules combined with neural architecture search and end-to-end optimization of the imputation pipeline.
    2. (b)
      auto-ml (Jäger et al., 2021): This approach uses the AutoML library autokeras (Jin et al., 2019) to implement the discriminative deep learning imputation method. For categorical columns, autokeras’ StructuredDataClassifier is employed, while StructuredDataRegressor is utilized for numerical columns. These classes manage data encoding and optimize model architecture and hyperparameters autonomously.
@@
    Multiple imputation (joint) using boostclean (Krishnan et al., 2017): This approach treats the error correction task as a statistical boosting problem where a set of weak learners are composed into a strong learner. To generate the weak learners, BoostClean iteratively selects a single imputation technique, applies it to a training set (with missingness), and fits a new model on the newly imputed training set.

### v4

@@ 5.3. Stability of the Predictive Model
Training set missingness. Figure 10 shows Label Stability of the Random Forest model on diabetes as a function of training set missingness rate. Under MCAR and MAR missingness, most MVI techniques with the exception of boostclean, deletion, and clustering show good stability (comparable to the model trained on clean data), and are generally insensitive to missingness rates. deletion and clustering are the least stable methods and show a monotonic decrease in stability with increase in missingness rate, with strongest effect under MNAR. miss-forest, auto-ml, and nomi are the most robust MVI techniques and preserve stability of the clean model under all missingness settings and error rates. This further highlights how imputation quality directly influences data uncertainty, ultimately impacting the overall uncertainty of the final model, as discussed in Section 4.3.
@@
Figure 51 in Appendix 50 shows the Label Stability of the Random Forest model on diabetes as a function of test missingness rate. We find that test missingness rate has little effect on model stability under all missingness settings, with the exception of clustering, which shows lower label stability at higher missingness rates, most pronounced in scenario S7 (MAR train, MNAR test).
6. Running Time
Table 4. Table 4: Training time (in seconds) of MVI techniques averaged across single- and multi-mechanism scenarios (S1-3, S10). Imputers are sorted by running time on the folk_emp dataset, and datasets are ordered by the number of rows. Dataset shapes reflect training sets with 30% rows with nulls. Values represent mean running times across seeds, with standard deviations.
Imputer diabetes (633, 17) german (700, 21) folk__\__inc (12000, 10) law__\__school (16638, 11) bank (32003, 13) heart (56000, 11) folk__\__emp (242112, 16)
median-dummy $0.013\pm 0.000$ $0.014\pm 0.001$ $0.021\pm 0.001$ $0.024\pm 0.001$ $0.027\pm 0.000$ $0.053\pm 0.001$ $0.773\pm 0.024$
median-mode $0.012\pm 0.000$ $0.014\pm 0.001$ $0.023\pm 0.001$ $0.025\pm 0.001$ $0.031\pm 0.001$ $0.067\pm 0.003$ $0.933\pm 0.019$
deletion $0.013\pm 0.000$ $0.013\pm 0.001$ $0.024\pm 0.001$ $0.025\pm 0.000$ $0.046\pm 0.001$ $0.087\pm 0.004$ $1\pm 0.049$
mnar__\__pvae $8\pm 0.705$ $14\pm 11$ $14\pm 1$ $22\pm 11$ $55\pm 30$ $42\pm 6$ $206\pm 7$
edit__\__gain $2\pm 0.119$ $2\pm 0.141$ $13\pm 0.221$ $21\pm 2$ $30\pm 1$ $62\pm 7$ $215\pm 4$
nomi $11\pm 3$ $14\pm 7$ $22\pm 2$ $22\pm 2$ $29\pm 1$ $38\pm 2$ $356\pm 20$
tdm $932\pm 74$ $1023\pm 11$ $1297\pm 22$ $1172\pm 92$ $1412\pm 34$ $1310\pm 42$ $1449\pm 31$
notmiwae $161\pm 101$ $217\pm 82$ $555\pm 2$ $804\pm 8$ $665\pm 284$ $1393\pm 535$ $2944\pm 725$
gain $261\pm 12$ $298\pm 5$ $1115\pm 38$ $1484\pm 30$ $1964\pm 36$ $2892\pm 48$ $6148\pm 178$
hivae $68\pm 0.784$ $95\pm 1$ $745\pm 15$ $1073\pm 11$ $2450\pm 130$ $3794\pm 129$ $7163\pm 317$
k__\__means__\__clustering $10\pm 0.402$ $13\pm 0.442$ $27\pm 0.640$ $323\pm 7$ $998\pm 49$ $1037\pm 64$ $7427\pm 778$
miss__\__forest $111\pm 17$ $244\pm 86$ $1758\pm 526$ $2530\pm 851$ $4307\pm 1310$ $6337\pm 1601$ $20358\pm 4934$
datawig $596\pm 185$ $277\pm 59$ $604\pm 46$ $2361\pm 492$ $5089\pm 651$ $7592\pm 854$ $31060\pm 3743$
automl $1953\pm 195$ $1805\pm 212$ $5559\pm 581$ $6803\pm 565$ $13893\pm 1687$ $19055\pm 2710$ $104476\pm 14743$
Table 4 presents the training time of MVI techniques for each dataset, averaged across all unique training scenarios (single- and multi-mechanism S1-S3, S10). Our time efficiency analysis approach aligns with prior work (Wang et al., 2024b; Miao et al., 2021), who also focused on training time as inference times are comparably fast across all techniques.
Our results reveal that statistical imputers deletion, median-mode, and median-dummy are the fastest, while still delivering competitive accuracy for larger datasets like heart and folk_emp, as shown in Figures 18 and 19. In contrast, miss-forest, datawig, and auto-ml exhibit the longest training times, with at least one of these methods achieving the highest imputation accuracy in most cases. Interestingly, auto-ml requires three times more training time than datawig, the second most computationally intensive technique. This difference is due to the auto-ML nature of auto-ml, which involves extensive hyperparameter and network architecture tuning.
Among non-statistical techniques, mnar-pvae, edit-gain, and nomi are the most efficient in terms of runtime. Notably, nomi provides competitive accuracy comparable to miss-forest, datawig, and auto-ml, making it the most optimal technique in terms of the trade-off between imputation accuracy and training time. A particularly noteworthy comparison is between gain and edit-gain. As explained in Appendix A.1, EDIT accelerates the training of parametric imputation models. Applying EDIT to gain results in a 28x speedup on the folk_emp dataset, with even greater improvements for smaller datasets, as shown in Table 4. Importantly, with this running time acceleration, edit-gain achieves accuracy comparable to gain on most datasets, as shown in Figures 18 and 19.
@@ 7. Summary of Experimental Findings
Do not drop your nulls! Building on evidence in the scientific community (Little and Rubin, 1987; Martínez-Plumed et al., 2019; Little et al., 2015; Joel and Doorsamy, 2022), our experiments show that deletion is the least effective strategy for model accuracy, fairness, and stability, especially in setups where each row contributes valuable information. Logically, deletion results in data loss, but its appropriateness depends on data quality rather than quantity. If rows are complete duplicates or contain erroneous values, deletion may be justified.
@@
There is conflicting evidence on the performance of multiple imputation (MI) (Joost R. van Ginkel and van der Voort, 2020), and our empirical findings are similarly mixed and somewhat unexpected. Feng (2023) and Le Morvan et al. (2021) argue that MI outperforms impute-then-classify approaches in predictive performance, while Graham (2009) and McNeish (2017) find MI effective even with limited data and small error rates. In contrast, we find that MI (specifically, boostclean) is only competitive on small datasets and is less stable than simpler MVI techniques. This likely stems from the complexity-stability trade-off (Domingos, 2000; Breiman, 2001): MI employs a more complex model class that minimizes empirical loss but exhibits greater prediction variance under small training set perturbations.
@@
Shadbahr et al. (2023) argue that distributional imputation quality metrics better predict model performance than discrepancy metrics. However, we find that neither reliably indicates downstream performance (F1), as strong learners can compensate for poor imputation quality. Moreover, imputation fairness metrics do not predict model fairness: perfect imputation fairness can still lead to model unfairness, while models trained with imputation-unfair MVI techniques can achieve good downstream fairness.
@@
Figure 11. Spearman correlation ($\rho$) between MVI technique, model type, and performance metrics (F1, fairness and stability) for different train missingness mechanisms (subplots). TPRD and TNRD values close to 0 are ideal (fair), so we compute correlations using $TPRD\_Reversed=1-|TPRD|$ and $TNRD\_Reversed=1-|TNRD|$. Supplementary plots are in Appendix E.
@@
Model performance (F1) is more affected by test missingness than training missingness. Fairness, however, is highly sensitive to both. Model stability is largely unaffected by test missingness, but most MVI techniques become more unstable with higher training missingness.
Existing MNAR-specific methods are insufficient.
MNAR is theoretically the hardest setting to model. Empirically, we find that MVI techniques perform poorly with MNAR during testing, even the two existing MNAR-specific methods (not-miwae and mnar-pvae). For instance, experimental results show that model F1 and stability are more sensitive to missingness rates under MNAR than under MCAR and MAR. This performance gap arises because these methods rely on assumptions that often fail in real-world settings and are evaluated on limited, mostly large datasets with a single scenario, lacking the diversity of our benchmark. Developing improved MNAR-specific MVI methods and establishing more comprehensive benchmarks for their evaluation are critical directions for future research.
@@
Interest in deep learning for tabular data has surged (Hwang and Song, 2023; Gorishniy et al., 2021; Borisov et al., 2022; Ye et al., 2024; Fayaz et al., 2022). Recall that Figures 3, 4, and 5 show only the best-performing models (by F1) for clarity. Tree-based methods like random forests and gradient-boosted trees outperform deep table-learning models like GANDALF (Joseph and Raj, 2022) on small datasets (diabetes, german), aligning with prior findings (Grinsztajn et al., 2022; Shwartz-Ziv and Armon, 2022). However, GANDALF excels on larger datasets (heart, folk-employment), highlighting the increasing effectiveness of deep learning across diverse data modalities.
@@
A key novelty of our study is the evaluation under mixed or multi-mechanism missingness—a more complex but realistic pattern likely encountered in practice. We find that performance trends vary between single- and multi-mechanism missingness. Figure 11 shows that model types like gandalf__\__clf and lgbm__\__clf are uncorrelated with performance metrics under single-mechanism missingness but become correlated under mixed missingness: F1 correlations are $\rho$=-0.24 for gandalf__\__clf and $\rho$=-0.37 for lgbm__\__clf, while fairness (TPRD) correlations are $\rho$=0.55 for gandalf__\__clf and $\rho$=-0.4 for lgbm__\__clf.
@@
Prior studies (Hasan et al., 2021; Lin and Tsai, 2020; Joel and Doorsamy, 2022; Shadbahr et al., 2023; García Laencina et al., 2010) have found no universally best missing value imputation technique for predictive performance. Our findings reinforce this, revealing trade-offs between F1, fairness, and stability that depend on the predictive model’s architecture. For instance, under multi-mechanism missingness (Figure 11), logistic regression (lr__\__clf) correlates positively with F1 ($\rho$=0.23) and stability ($\rho$=0.38) but negatively with fairness (TPRD, $\rho$=-0.23). Random forest (rf__\__clf) shows weak positive correlations with fairness (TPRD, $\rho$=0.16) and F1 ($\rho$=0.25) but a negative correlation with stability ($\rho$=-0.24). Deep table-learning (gandalf__\__clf) is strongly correlated with fairness (TPRD, $\rho$=0.55), weakly with stability ($\rho$=0.12), and negatively with F1 ($\rho$=-0.24). This is not bad news, but rather an acknowledgement of the complexity of learning from incomplete data and the need for rigorous, holistic evaluation protocols, such as those demonstrated in this work, to select the most suitable imputation method and model architecture for a given task.
@@ 8. Conclusions, Limitations and Future Work
Conclusions. We introduced Shades-of-Null, an evaluation suite for responsible missing value imputation. A key contribution is evaluating fairness and stability alongside predictive performance. Additionally, we model realistic missingness scenarios beyond Rubin’s MCAR, MAR, and MNAR, incorporating multi-mechanism missingness and missingness shift. Through 29,736 experimental pipelines, we assessed various MVI methods under realistic missingness conditions, revealing key trends and trade-offs across evaluation metrics.
@@
Creating research datasets with naturally occurring missingness shifts is a promising direction. While predictive performance, fairness, and stability are often treated as orthogonal, we uncover trade-offs driven by missingness, imputation choice, and, in some cases, model architecture. The absence of a universally best imputation method highlights the need for holistic metrics, evaluation procedures, and techniques grounded in the data lifecycle. Missing value imputation remains a critical aspect of responsible data engineering, to which our work makes a contribution.
@@
This supplemental material provides significantly more detail regarding the results presented in this paper. The organization is as follows: We first provide additional details on the implementation of the evaluation suite, which include a broad description and configuration of the MVI techniques used in our study (Appendix A.1) and an explanation of the corrections and enhancements made to existing benchmarks and MVI techniques to ensure the development of our truly fair evaluation suite (Appendix A.2). Next, we extend the description of our experimental settings to facilitate reproducibility (Appendix B). Specifically, Appendix B.1 provides detailed information about our experimental setup on the cluster, while Appendix B.2 extends our methodology for simulating missingness in other datasets, including proportions and base rates of protected groups, realistic missingness scenarios, feature correlation coefficients with the target, and feature importance for each dataset. In Appendix C, we offer additional results to those discussed in Section 4. In particular, Appendix C.1 presents the imputation quality of MVI techniques across all datasets from the perspectives of accuracy and training time, while Appendix C.2 showcases the fairness of predictive models measured by metrics beyond the TPRD mentioned in Section 4 and includes scatter plots for all imputers that extend scatter plots from Section 4.4. Then, Appendix D presents supplementary experimental findings on the missingness shift discussed in Section 5. Specifically, Appendix D.1 examines the accuracy, stability, and fairness dimensions of model performance under missingness shift with fixed train and test error rates, highlighting its effects on various model types and datasets. Furthermore, Appendix D.2 explores further experimental results on model performance under missingness shift with variable train and test error rates. Finally, Appendix E presents correlation plots for other missingness settings than in Section 7.
@@ A.1. Missing Value Imputation Techniques
In our evaluation suite, we evaluate 15 MVI techniques, classified into 8 broad categories, as outlined in Section 3.2. These categories comprehensively cover the spectrum of MVI methods discussed in recent surveys (Emmanuel et al., 2021; Jäger et al., 2021; Miao et al., 2022; Abdelaal et al., 2023) and peer-reviewed papers (Wang et al., 2024b; Wu et al., 2023; Miao et al., 2021), which propose new MVI techniques and compare against them. In each category, we selected several SOTA methods for our comparison. Each technique is capable of imputing both categorical and numerical features. For all MVI techniques, we adopt the default parameter settings from original papers or their source code.
@@
    median-mode (statistical): We impute missing values with the median of the complete case for numerical columns and mode (most frequently occurring value) for categorical columns.
@@
    median-dummy (statistical): We impute missing values with the median of the complete case for numerical columns and assign a (new) dummy category for nulls in categorical columns.
@@
    miss-forest (Stekhoven and Bühlmann, 2011) (supervised ML): This approach iteratively trains an RF model on a set of clean samples (no missingness) and predicts the missing values. We tune RandomForestClassifier for categorical columns with nulls and RandomForestRegressor for numerical columns with nulls. The maximum number of iterations for miss-forest is set to 10.
@@
    clustering (Gajawada and Toshniwal, 2012) (unsupervised ML): We assign the missing value to a cluster, based on the distance from the cluster center, and then impute the missing value with the mean or mode of the data points in that cluster. [^2] We used the k-prototypes implementation from https://github.com/nicodv/kmodes We implemented this method ourselves, as no existing solution supports null imputation based on clustering for data with both numerical and categorical features.
  6. (6)
    datawig (Biessmann et al., 2019) (discriminative DL): This method implements deep learning modules combined with neural architecture search and end-to-end optimization of the imputation pipeline. We train datawig for 100 epochs and a single iteration using a batch size of 64. The final_fc_hidden_units parameter is tuned for each dataset by selecting the optimal value from the set {1, 10, 50, 100}.
  7. (7)
    auto-ml (Jäger et al., 2021) (discriminative DL): This approach uses the AutoML library autokeras (Jin et al., 2019) to implement the discriminative deep learning imputation method. For categorical columns, autokeras’ StructuredDataClassifier is employed, while StructuredDataRegressor is utilized for numerical columns. These classes manage data encoding and optimize model architecture and hyperparameters autonomously. We train auto-ml for 100 epochs with a validation split of 0.2, using up to 50 trials to tune hyperparameters.
  8. (8)
    gain (Yoon et al., 2018) (generative DL): This method adapts the Generative Adversarial Networks (GAN) framework to impute missing data. The generator observes partially observed data and imputes the missing components conditioned on the observed data, outputting a completed data vector. The discriminator distinguishes between observed and imputed components, using a hint vector to focus on specific imputation tasks. The adversarial training ensures that the generator learns to produce imputations that closely match the true data distribution. We train gain by tuning its hyperparameters based on the following grid: $\alpha\in\{1,10\}$, hint rate $\in\{0.7,0.9\}$, batch size $\in\{64,128\}$, generator learning rate $\in\{1\times 10^{-5},1\times 10^{-4},0.0005\}$, and discriminator learning rate $\in\{1\times 10^{-6},1\times 10^{-5},0.00005\}$.
  9. (9)
    hi-vae (Nazabal et al., 2020) (generative DL): This approach extends the Variational Autoencoder (VAE) framework to handle incomplete and heterogeneous datasets. HI-VAE models mixed numerical (e.g., real-valued, positive real-valued, and count) and categorical (e.g., ordinal and nominal) data types. It employs a hierarchical architecture with a shared latent space to capture correlations among attributes and a likelihood model tailored to each data type. By training on observed data and utilizing an Evidence Lower Bound (ELBO) computed only for observed entries, HI-VAE achieves robust imputation without overfitting. We train hi-vae for 2000 epochs using a batch size of 128, with latent dimensions set as $z=10$, $y=5$, and $s=10$, and a learning rate of $1\times 10^{-3}$.
  10. (10)
    not-miwae (Ipsen et al., 2020) (MNAR-specific): This method extends deep generative modeling to handle Missing Not At Random (MNAR) data by explicitly modeling the missing data mechanism alongside the data distribution. It leverages importance-weighted variational inference to jointly optimize the parameters of the data and missing mechanisms. The method uses stochastic gradients derived via reparameterization in both latent and data spaces, allowing efficient training. By incorporating prior knowledge about the missingness process into a flexible missing model, not-miwae achieves robust imputation under MNAR scenarios. We train not-miwae for a maximum of 100,000 iterations using a batch size of 16, with 128 hidden units, and a latent dimensionality of $L=10,000$, employing the self-masking process selfmasking_known.
  11. (11)
    mnar-pvae (Ma and Zhang, 2021) (MNAR-specific): This approach introduces an identifiable deep generative model for handling Missing Not At Random (MNAR) data. By leveraging identifiability principles and extending variational autoencoders (VAEs), GINA ensures that the underlying data-generating process can be uniquely recovered. It jointly models the data distribution and the missingness mechanism, integrating auxiliary variables to achieve identifiability under mild assumptions. Through a combination of importance-weighted variational inference and a flexible neural architecture, GINA enables robust and unbiased imputation of MNAR data. We train mnar-pvae (GINA) for 400 epochs and a single iteration using a batch size of 100 and a learning rate of $1\times 10^{-3}$. The network is configured with an embedding dimension of 20, a latent dimension of 20, encoder and decoder layers each containing 10 units, and uses Tanh as the non-linearity.
  12. (12)
@@
    Multiple imputation (joint) using boostclean (Krishnan et al., 2017): This approach treats the error correction task as a statistical boosting problem where a set of weak learners are composed into a strong learner. To generate the weak learners, BoostClean iteratively selects a single imputation technique, applies it to a training set (with missingness), and fits a new model on the newly imputed training set. We train boostclean for 5 iterations and tune its underlying prediction model for each iteration.
  13. (13)
    nomi (Wang et al., 2024b) (recent): This method introduces an uncertainty-driven network for missing data imputation. NOMI integrates three key components: a retrieval module, a Neural Network Gaussian Process Imputator (NNGPI), and an uncertainty-based calibration module. The retrieval module identifies local neighbors of incomplete samples, while the NNGPI combines the probabilistic modeling of Gaussian Processes with the representation power of neural networks to impute missing values and quantify uncertainty. The uncertainty-based calibration module dynamically refines the imputations by balancing predictions across iterations, leveraging uncertainty to improve reliability. We train nomi with a maximum of 3 iterations, using 10 neighbors for imputation, the $l_{2}$ similarity metric, a temperature parameter $\tau=1.0$, and a weighting coefficient $\beta=0.8$.
  14. (14)
    tdm (Zhao et al., 2023) (recent): This approach proposes Transformed Distribution Matching (TDM), a novel approach to missing data imputation that leverages optimal transport in a transformed latent space. TDM employs deep invertible neural networks to map data into a latent space where the geometry better reflects the underlying data structure. It minimizes the Wasserstein distance between the transformed distributions of two batches of data to achieve imputation, while avoiding overfitting through mutual information constraints. We train tdm for 10,000 iterations with a batch size of 512 and a learning rate of $1\times 10^{-2}$.
  15. (15)
    edit-gain (Miao et al., 2021) (recent): This technique introduces an efficient and effective data imputation framework that leverages influence functions to accelerate the training of parametric imputation models. EDIT consists of two key modules: the Imputation Influence Evaluation (IIE) module, which estimates the influence power of samples on the imputation model’s predictions, and the Representative Sample Selection (RSS) module, which constructs a minimal representative sample set to satisfy user-specified accuracy guarantees. Additionally, EDIT employs a weighted loss function that emphasizes high-influence samples, boosting imputation accuracy while reducing training cost. We employ gain (Yoon et al., 2018) as the underlying imputer for EDIT, as it is the only method shared with us directly by the authors of the original paper. We train edit-gain using a batch size of 128 for 30 epochs, with an $\alpha$ parameter value of 1. The initial sample size is set to 6000 for all our datasets, matching the initial sample size of the smallest dataset used in the original paper, since all our datasets are of the same size or smaller.